Please help me find Volts and the Amps on each resistor.
Circuit 5 A Xe R1 25 Ohms R2 30 Ohms A- 60 VDC R4 60 Ohms R3 30 Ohms Rts R6 50 Ohms 10 Ohms
Answer #1

RL 60 3+32 = 60-70.93 3S 1.116 A C TSD SD 60 60 30430 2o-73I-11C 53·46-а VR4 = 20.93V 2